# HeadGaS: Real-Time Animatable Head Avatars via 3D Gaussian Splatting

### v1

@@ Abstract
3D head animation has seen major quality and runtime improvements over the last few years, particularly empowered by the advances in differentiable rendering and neural radiance fields. Real-time rendering is a highly desirable goal for real-world applications. We propose HeadGaS, the first model to use 3D Gaussian Splats (3DGS) for 3D head reconstruction and animation. In this paper we introduce a hybrid model that extends the explicit representation from 3DGS with a base of learnable latent features, which can be linearly blended with low-dimensional parameters from parametric head models to obtain expression-dependent final color and opacity values. We demonstrate that HeadGaS delivers state-of-the-art results in real-time inference frame rates, which surpasses baselines by up to $2dB$, while accelerating rendering speed by over ×$\times 10$.
@@ 1 Introduction
Animatable 3D head reconstruction consists on driving a captured head avatar, based on a target sequence of facial expressions and head poses. In the last decades, various parametric 3D morphable models (3DMM) have emerged [4, 20, 5], which can be fitted to sequences of a moving head and later on enable pose and expression control. Though these models make it possible to drive a captured avatar via a set of low-dimensional parameters, generally their generated images lack realism. Other works utilize the fitting of low-dimensional parameters from such 3DMM models for initial estimates and build on other mechanisms to obtain more realistic imagery with animation capabilities [13, 10].
Figure 1: Overview of HeadGaS. We reconstruct a 3D head based on an expression-aware 3D Gaussian cloud representation, which results in real-time rendering and high image quality. Left: The model is trained with a monocular video of a moving head. Right: At inference, we query the model with a novel sequence of camera poses and expression parameters to render a real-time video.
@@
In particular, with the recent success of differentiable rendering, various 3D-aware animatable head models emerged that can reconstruct and render 3D heads, while providing the functionality to drive them based on expression parameters from 3DMM models. These representations can be explicit (mesh, point clouds) [13, 47] or implicit (neural) [11]. Thereby, the explicit models impose stronger constraints on the head surface, which allows for better expression and pose generalization, while making it more difficult to preserve photo realism, as they inherit the limitations and artifacts of the underlying representation (mesh, point cloud) as reported in findings from other works [11] and also seen in our experiments (Fig. 3). With the recent success of neural radiance fields (NeRFs) [24], typically implicit models are based on a NeRF representation [10]. Some of these models [48, 11] prioritize time constraints and therefore rely on very fast volumetric NeRF variants (e.g. Instant NGP [25]) to enable fast training and rendering.
@@
Despite impressive efforts to improve NeRFs to be more accurate [3] and fast [25], there is a trade-off between these two aspects that is hard to satisfy simultaneously. Moreover, even fast and efficient NeRF models like InstantNGP typically enable interactive inference frame rates at best (10-15 fps) [17]. Very recently, 3D Gaussian Splatting (3DGS) [17] emerged as a competitive alternative to NeRF, which leads to reasonable photo-realism while bringing the rendering speed to real-time rates. This is thanks to its representation as a set of 3D Gaussian primitives, with a more efficient space coverage compared to point clouds, combined with efficient tile-based rasterization. However, in the light of 3D head animation, in its original form, 3DGS does not constitute an intuitive surface or point set that can be directly deformed based on 3DMM deformation, unlike other well-known representations e.g. surface or pointcloud based.
@@
To circumvent this limitation, we propose HeadGaS, the first work that enhances 3D Gaussians with head animation capabilities (see Figure 1). At test time, our model receives a sequence of camera views and expression parameters and is capable of generating a corresponding video of the reconstructed avatar. Thereby, inspired by the expression blending idea of traditional 3DMMs [4], we introduce a base of latent features inside each Gaussian. This base is multiplied with an expression vector, and its sum is fed to a multi-layer perceptron (MLP) to yield the final color and opacity. Such model allows for varying colors of a certain rendered frame based on the expression vector. Our model can work with any 3DMM representation, as it does not explicitly model deformations with respect to a particular mesh topology. Practically, in our experiments we show that HeadGaS can be controlled with expression parameters from two different 3DMMs, namely FLAME [20] and FaceWarehouse [5]. The rendering is done in real-time frame rates, at over 100fps (more than $200$fps for $512^{2}$ resolution).
@@
We evaluate our model in publicly monocular video datasets, also used in related works [11, 48, 47]. Thereby, we demonstrate that the proposed model is capable of producing state-of-the-art results, while increasing the rendering speed by a factor of at least ×$\times 10$ compared to interactive NeRF-based baselines [11, 48]. We show the applications of HeadGaS in various tasks, such as novel same-person expression transfer, cross-subject expression transfer, as well as novel view synthesis.
@@
    We formulate the first work that can render animatable heads on real-time, adopting an efficient set of 3D Gaussian primitives as representation.
@@
    We extend the 3D Gaussian representation [17] with a base of latent features, which can be weighted by an expression vector to enable head expression controllability.
@@
    We evaluate extensively our proposed method, and compare it against other recent state-of-the-art approaches, obtaining up to 2dB improvement and ×$\times 10$ speed ups.
@@ 2 Related Work
Our work consists in reconstruction and rendering of controllable 3D heads, via a set of images. As our framework relies on a radiance field, we first give an overview of the respective relevant methods. Further, we discuss works related to animatable head reconstruction.
2.1 Radiance Fields
NeRF [24] represent the scene as an implicit neural radiance field, that queries 3D space and predicts density and view-dependent color via a multi-layer perceptron (MLP). In the following years, many follow-up works have focused on improving different aspects of it such as anti-aliasing  [1, 2, 3], regularization for sparse views [26, 8, 37] and speed  [35, 25, 6]. DVGO [35] replace the MLP of NeRF with a density and learned feature voxel grid to considerably speed up convergence. TensoRF [6] factorize the 4D feature voxel grid of a scene into a set of low-rank 2D and 3D tensors which improves efficiency. InstantNGP [25] employ a hash grid and an occupancy grid to accelerate computation, followed by a small MLP that infers density and color. NeRFs have also been used to represent dynamic scenes including human bodies [39, 7, 30], human heads [11, 48], and generic time-varying scenes [31, 29, 21, 9, 28, 36]. Typically these models rely on a canonical space, where all observations are mapped for time consistent reconstruction. Thereby, the fastest head animation methods [48, 11] build on an InstantNGP hash grid to enable rendering in interactive frame rates (10-15 fps).
3DGS [17] represent a scene as a set of explicit 3D Gaussians with the motivation to minimize computation in empty spaces. Their efficient representation combined with tile-based rasterization algorithm allows for accelerated training and real-time rendering (over 100fps). A line of works extends the 3D Gaussian representation to model dynamic scenes [22, 40, 43]. Luiten et al. [22] propose simultaneous dynamic scene reconstruction and 6-DoF tracking by allowing the Gaussians to change position and rotation over time while enforcing the same color, size and opacity. Yang et al. [43] learn a MLP based deformation that maps 3D Gaussians to a canonical space. 4D-GS [40] propose an efficient deformation field by querying features in shared multi-resolution voxel planes. To the best of our knowledge, we are the first to propose a 3D head animation solution, and extend the explicit 3DGS representation with neural features for blending.
Figure 2: HeadGaS pipeline. We represent 3D space as a set of feature-enhanced 3D Gaussians. Every Gaussian contains a feature basis $\bm{F}$ that can be blended via the expression vector to obtain a frame specific feature $\bm{f}_{i}$. The frame specific feature is fed to an MLP ϕ()italic-$\phi(\cdot)$ alongside position $\bm{\mu}$ to obtain expression-dependent color $\bm{c}_{i}$ and opacity $\bm{\alpha}_{i}$. Finally, $\bm{c}_{i}$ and $\bm{\alpha}_{i}$ are fed to the rasterizer alongside other Gaussian parameters like rotation $R$, scale $S$ and position $\bm{\mu}$ to render the image.
@@ 2.2 Head reconstruction and animation
Head reconstruction from a set of image observations has been a very active field in the recent years, including models that generalize across subjects  [23, 14, 38], or rely on multi-view head captures [19, 15], which can have a static [38, 15] or dynamic form [14, 19].
Most related to our proposed method, are works that learn a dynamic, animatable 3D head model from a monocular video sequence that observes the head in various poses and facial expressions, and is capable to generate a novel expression or pose at test time. A line of works rely on explicit scene representations, such as meshes or point clouds [18, 12, 13, 47]. Neural Head Avatars [13] models the geometry and texture explicitly via a hybrid representation consisting of a coarse morphable model, followed by a neural based refinement that predict voxel offsets. PointAvatar [47] propose a deformable point-based representation, where all points have the same radius. Our 3D Gaussian set shares some similarity with a point cloud, yet it is more flexible. Each Gaussian can have its own radius and different axis lengths (i.e. ellipsoidal shape instead of spherical).
Another line of works extend implicit neural radiance representations. NerFACE [10] use a dynamic NeRF to combine scene information with a morphable head model to enable pose and expression control. IMAvatar [46] utilizes neural implicit surfaces [27] and learns an implicit deformation field from canonical space to observation based on expression parameters and pose. With the goal of fast training, and interactive rendering, more recent works extend IntantNGP [25] with head aware models. INSTA [48] use a tracked FLAME mesh as a geometrical prior to deform the point into canonical space, followed by InstantNGP [25]. NeRFBlendShape [11] follow a different approach, that in contrast to most previous works does not rely on deformation. Instead, inspired by classic blendshape models for heads [5], they utilize a base of multi-level hash grid fields [25], where the model can be driven via a linear blending of such hash grid base with the expression vector. Similarly AvatarMav [42] use expression weights to blend a set of motion voxel grids. Our proposed approach resembles these blending-based ideas, but instead we use the expression vector to blend latent Gaussian features to predict expression-specific color and opacity.
Different from all works discussed in this section, we adopt a set of 3D Gaussian as neural radiance representation, to take advantage of the fast rendering benefits, combined with competitive photo realism.
@@ 3 Method
Given a monocular video of a moving head, our goal is to learn a 3D head model and render novel images of this avatar based on a facial expression vector and camera pose. As a pre-processing step, similar to other works [11, 48, 13, 47], we require head poses and a vector of expression weights associated with each frame, and adopt a head tracking pipeline to achieve this. Note that HeadGaS does not explicitly build on a certain parametric model, and it can therefore work with different head models. Practically we have performed experiments with [20, 5] based on the head tracking frameworks of  [11, 48, 46]. The resulting head poses are converted to camera poses. In addition, we perform video matting followed by face parsing to discard background areas and clothing to focus on the head area only. In the following sections, we will first provide some background on the vanilla 3DGS (Sec. 3.1) and further describe the proposed strategy for animatable 3DGS (Sec. 3.2), the rendering (Sec. 3.3) and its optimization (Sec. 3.4).
3.1 Vanilla 3DGS Representation
Given a set of images of a static scene and the corresponding camera poses, 3DGS [17] learn a 3D scene representation as a set of 3D gaussians, and can render a novel image from a given viewpoint. To initialize the 3D Gaussians, 3DGS utilizes a sparse point cloud, typically originating from the COLMAP [33] framework by which they also obtain the camera poses. Thereby, a 3D Gaussian is represented as a tuple of 3D covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$, Gaussian center $\bm{\mu}\in\mathbb{R}^{3}$, color $\bm{c}\in\mathbb{R}^{3(k+1)^{2}}$ and opacity $\alpha\in\mathbb{R}$, i.e. $\mathcal{G}=(\bm{\Sigma},\bm{\mu},\bm{c},\alpha)$, where $k$ is the degree of the spherical harmonics. The Gaussians are defined in world space, centered at the mean point as
@@ 3.1 Vanilla 3DGS Representation
To make optimization stable, i.e. guarantee that $\bm{\Sigma}$ is positive semi-definite, the covariance matrix is further decomposed into rotation and scaling:
@@
Color is represented via spherical harmonics (SH) with degree $k$ and is therefore view-dependent.
The Gaussian parameters are optimized via a differentiable rasterizer, that projects the current 3D Gaussians to the image space and compares against the given ground truth images. The differentiable rasterizer relies on an efficient algorithm for sorting the Gaussians and tiling the image space, which leads to very fast training and rendering. Alongside the optimizations, 3DGS employs an adaptive mechanism for pruning and densification, to make sure that the set of gaussians represents the space effectively. For more details, we refer the reader to the 3DGS paper [17].

### v2

@@ Abstract
3D head animation has seen major quality and runtime improvements over the last few years, particularly empowered by the advances in differentiable rendering and neural radiance fields. Real-time rendering is a highly desirable goal for real-world applications. We propose HeadGaS, a model that uses 3D Gaussian Splats (3DGS) for 3D head reconstruction and animation. In this paper we introduce a hybrid model that extends the explicit 3DGS representation with a base of learnable latent features, which can be linearly blended with low-dimensional parameters from parametric head models to obtain expression-dependent color and opacity values. We demonstrate that HeadGaS delivers state-of-the-art results in real-time inference frame rates, surpassing baselines by up to $2$ dB, while accelerating rendering speed by over ×$\times 10$.
Keywords:
animatable head avatars gaussian splatting radiance fields
Figure 1: Overview of HeadGaS. We reconstruct a 3D head based on an expression-aware 3D Gaussian cloud representation, which results in real-time rendering and high image quality. Left: The model is trained with a monocular video of a moving head. At inference, we query the model with a novel sequence of poses and expression parameters to render a real-time video. Right: Rendering speed (fps in logarithmic scale) vs PSNR plot comparing different methods. The circle radius indicates training time.
@@ 1 Introduction
Animatable 3D head reconstruction consists in driving a captured head avatar, based on a target sequence of facial expressions and head poses. In the last decades, various parametric 3D morphable models (3DMM) have emerged [5, 24, 6], which can be fitted to sequences of a moving head and later on enable pose and expression control. Though these models make it possible to drive a captured avatar via a set of low-dimensional parameters, generally their generated images lack realism. Other works utilize the fitting of low-dimensional parameters from such 3DMM models for initial estimates and build on other mechanisms to obtain more realistic imagery with animation capabilities [16, 13].
@@
In particular, with the recent success of differentiable rendering, various 3D-aware animatable head models emerged that can reconstruct and render 3D heads, while providing the functionality to drive them based on expression parameters from 3DMM models. These representations can be explicit (mesh, point clouds) [16, 61] or implicit (neural) [14]. Thereby, the explicit models impose stronger constraints on the head surface, which allows for better expression and pose generalization, while making it more difficult to preserve photo realism, as they inherit the limitations and artifacts of the underlying representation (mesh, point cloud) as observed in Gao et al. [14] and also seen in our experiments (Fig. 4). With the recent success of neural radiance fields (NeRFs) [31], typically implicit models are based on a NeRF representation [13]. Some of these models [62, 14] prioritize time constraints and therefore rely on very fast volumetric NeRF variants (e.g. InstantNGP [33]) to enable fast training and rendering.
@@
Despite impressive efforts to improve NeRFs to be more accurate [3] and fast [33], there is a trade-off between these two aspects that is hard to satisfy simultaneously [3]. Moreover, even fast and efficient NeRF models like InstantNGP typically enable interactive inference frame rates at best (10-15 fps) [20]. Very recently, 3D Gaussian Splatting (3DGS) [20] emerged as a competitive alternative to NeRF, which leads to reasonable photo-realism while bringing the rendering speed to real-time rates. This is thanks to its representation as a set of 3D Gaussian primitives, with a more efficient space coverage compared to point clouds, combined with efficient tile-based rasterization. However, in the light of 3D head animation, in its original form, 3DGS does not constitute an intuitive surface or point set that can be directly deformed based on 3DMM deformation, unlike other well-known representations, e.g. surface or pointcloud based.
Figure 2: Motion modelling via opacity change. Left: Two example frames $i$ and $j$ rendered by HeadGaS. Right: Rendering of opacity difference $\alpha_{j}-\alpha_{i}$ (blue: Gaussians with an opacity decrease; red: Gaussians with an opacity increase; colors close to white: minor change, static regions). We observe a strong opacity increase in dynamic areas, e.g. lower chin Gaussians turn opaque as the jaw fully opens.
@@
To circumvent this limitation, we propose HeadGaS, a model that enhances 3D Gaussians with head animation capabilities (see Figure 1). At test time, our model receives a sequence of head poses and expression parameters, and generates a photo-realistic video of the reconstructed avatar. The core idea behind HeadGaS is to represent motion by allowing Gaussians to alter their opacity and color over time. As a consequence, HeadGaS will result in duplicates of the dynamic face areas, i.e. achieving dynamics via over-representation. Thus, there will be multiple Gaussians representing the same face region, and these duplicates will become active (apparent) one at a time, to support the face geometry at a certain state of expression. Figure 2 illustrates the opacity change as a result on an expression transition. Note how Gaussians representing lower lip and chin areas in frame $i$ turn transparent to allow for mouth opening in frame $j$, while another set of chin Gaussians emerge at a lower location to accommodate the new jaw position in frame $j$. To allow for such varying appearance, guided by an expression vector, we introduce a basis of latent features inside each Gaussian. This learned basis is multiplied with an input expression vector, and its sum is fed to a multi-layer perceptron (MLP) to yield the final color and opacity. This idea is inspired by traditional blendshape 3DMMs [6], and it can be interpreted as a latent-feature shape basis that is blended in the feature domain rather than directly in PCA space [5] or meshes [6]. Our model is simple and effective, and it can work with various 3DMM representations, as it does not explicitly model deformations with respect to a particular mesh topology. Practically, in our experiments we show that HeadGaS can be controlled with expression parameters from two different 3DMMs, namely FLAME [24] and FaceWarehouse [6]. The rendering is done in real-time framerates, at over 100 fps (about $250$ fps for $512^{2}$ resolution). We show experimentally that our visibility-varying Gaussians outperform the evident alternative of moving the Gaussians, which we attribute to the fact that adding 3D motion makes the optimization even more complex.
@@
We evaluate our model on publicly available monocular video datasets, commonly used in related works [14, 62, 61]. Thereby, we demonstrate that the proposed model yields superior results, while increasing the rendering speed by at least a ×$\times 10$ factor compared to interactive NeRF-based baselines [14, 62]. We show the applications of HeadGaS in novel same-person expression transfer, cross-subject expression transfer, as well as novel view synthesis.
@@
    We formulate a novel framework that can render photo-realistic 3D-aware animatable heads in real-time, adapting an efficient set of 3D Gaussian primitives. This framework handles face dynamics by allowing opacity and color to change over time, i.e. leveraging over-representation.
@@
    We extend 3DGS [20] with a per-Gaussian basis of latent features, which can be blended with expression weights to enable expression control.
@@
    We extensively evaluate our proposed method, and compare it against state-of-the-art approaches, obtaining up to 2dB improvement and ×$\times 10$ speed-ups.
@@ 2 Related Work
2.1 Towards Fast and Dynamic Radiance Fields
NeRF [31] represent the scene as an implicit neural radiance field, that queries 3D space and predicts density and view-dependent color via a MLP. In the following years, many follow-up works have focused on improving different aspects of it such as anti-aliasing  [1, 2, 3], regularization for sparse views [34, 11, 47] and speed  [45, 33, 8]; or enhancing results post-rendering [7]. DVGO [45] replace the MLP of NeRF with a density and learned feature voxel grid to considerably speed up convergence. TensoRF [8] factorize the 4D feature voxel grid of a scene into a set of low-rank 2D and 3D tensors which improves efficiency. InstantNGP [33] employ a hash grid and an occupancy grid to accelerate computation, followed by a small MLP that infers density and color. NeRFs have also been used to represent dynamic scenes including human bodies [50, 9, 38], human heads [14, 62], and generic time-varying scenes [39, 37, 25, 12, 36, 46]. Typically these models rely on a canonical space, where all observations are mapped for time consistent reconstruction. Methods aiming at fast rendering speeds  [62, 14] build on an InstantNGP hash grid and achieve interactive frame rates (10-15 fps).
3DGS [20] represent a scene as a set of explicit 3D Gaussians with the motivation to minimize computation in empty spaces. Their efficient representation combined with tile-based rasterization algorithm allows for accelerated training and real-time rendering (over 100 fps). A line of works extends the 3D Gaussian representation to model dynamic scenes [29, 51, 56, 43]. Luiten et al. [29] propose simultaneous dynamic scene reconstruction and 6-DoF tracking by allowing the Gaussians to change position and rotation over time while enforcing the same color, size and opacity. Yang et al. [56] learn a MLP based deformation that maps 3D Gaussians to a canonical space. 4D-GS [51] propose an efficient deformation field by querying features in shared multi-resolution voxel planes.
Very recent concurrent works model human head [54, 52, 10] and body [32, 23] avatars with 3DGS, by deforming a canonical head via a 3DMM-conditioned MLP [54, 52], relying on a tri-plane [49], or binding 3D Gaussians in a FLAME mesh [40]. Our method is quite different from these works, in that we utilize a per-Gaussian feature basis and opacity induced dynamics.
@@ 2.2 Head Reconstruction and Animation
Head reconstruction from a set of image observations has been a very active field in the recent years, including models that generalize across subjects [30, 17, 48], or rely on multi-view head captures [22, 18, 27, 28], which can have a static [48, 18] or dynamic form [17, 22]. Most related to our proposed method, are works that learn a dynamic, animatable 3D head model from a monocular video sequence that observes the head in various poses and facial expressions, and is capable to generate a novel expression or pose at test time. A line of works rely on explicit scene representations, such as meshes or point clouds [21, 15, 16, 61]. Neural Head Avatars [16] models the geometry and texture explicitly via a hybrid representation consisting of a coarse morphable model, followed by a neural based refinement that predict voxel offsets. PointAvatar [61] propose a deformable point-based representation, where all points have the same radius. Our 3D Gaussian set shares some similarity with a point cloud, yet it is more flexible. Each Gaussian can have its own radius, orientation and different axis lengths. Recently, FLARE [4] was proposed, a model that updates a traditional graphics pipeline with a few neural components. FLARE can optimize a 3D mesh via differentiable rendering, enabling avatars that are animatable and relightable.
Another line of works extend implicit neural radiance representations. NerFACE [13] use a dynamic NeRF to combine scene information with a morphable head model to enable pose and expression control. IMAvatar [60] utilizes neural implicit surfaces [35] and learns an implicit deformation field from canonical space to observation based on expression parameters and pose. With the goal of fast training, and interactive rendering, more recent works extend InstantNGP [33] with head aware models. INSTA [62] use a tracked FLAME mesh as a geometrical prior to deform the points into canonical space, followed by InstantNGP [33]. NeRFBlendShape [14] follow a different approach, that in contrast to most previous works does not rely on deformation. Instead, inspired by classic blendshape models for heads [6], they utilize a base of multi-level hash grid fields [33], where the model can be driven via a linear blending of such hash grid base with the expression vector. Similarly, AvatarMAV [55] use expression weights to blend a set of motion voxel grids. Our proposed approach resembles these blending-based ideas, but instead we use the expression vector to blend latent per-Gaussian features to predict expression-specific color and opacity.
Different from all works discussed in this section, we adopt a set of 3D Gaussians as neural radiance representation, to take advantage of the fast rendering benefits, combined with competitive photorealism.
Figure 3: HeadGaS pipeline. We represent 3D space as a set of feature-enhanced 3D Gaussians. Every Gaussian contains a feature basis $\bm{F}$ that can be blended via the expression vector to obtain a frame specific feature $\bm{f}_{i}$. The frame specific feature is fed to an MLP $\phi(\cdot)$ alongside position $\bm{\mu}$ to obtain expression-dependent color $\bm{c}_{i}$ and opacity $\bm{\alpha}_{i}$. Finally, $\bm{c}_{i}$ and $\bm{\alpha}_{i}$ are fed to the rasterizer alongside other Gaussian parameters like rotation $R$, scale $S$ and position $\bm{\mu}$ to render the image.
@@ 3 Method
Given a monocular video of a moving head, our goal is to learn a 3D head model and render novel images of this avatar based on a facial expression vector and camera pose. As a pre-processing step, similar to other works [14, 62, 16, 61], we require head poses and a vector of expression weights associated with each frame, and adopt a head tracking pipeline to achieve this. Note that HeadGaS does not explicitly build on a certain parametric model, and it can therefore work with different head models. We have performed experiments with two models [24, 6] based on the head tracking frameworks of prior work [14, 62, 60]. The resulting rigid head poses are converted to camera poses, to map all observations to the canonical head pose. In addition, we perform video matting [26] followed by face parsing [58] to discard background areas and clothing to focus on the head region only. In the following sections, we will first provide a background on the original 3DGS (Sec. 3.1) and further describe the proposed strategy for animatable 3DGS (Sec. 3.2), the rendering (Sec. 3.3) and its optimization (Sec. 3.4).
3.1 Original 3DGS Representation
Given a set of images of a static scene and the corresponding camera poses, 3DGS [20] learn a 3D scene as a set of 3D Gaussians, and can render a novel image from a given viewpoint. For initialization, 3DGS utilizes a sparse point cloud, typically originating from the COLMAP [42] framework by which they also obtain the camera poses. Thereby, a 3D Gaussian is represented as a tuple of 3D covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$, Gaussian center $\bm{\mu}\in\mathbb{R}^{3}$, color $\bm{c}\in\mathbb{R}^{3(k+1)^{2}}$ and opacity $\alpha\in\mathbb{R}$, i.e. $\mathcal{G}=(\bm{\Sigma},\bm{\mu},\bm{c},\alpha)$, where $k$ is the degree of the spherical harmonics. The Gaussians are defined in world space, centered at the mean point
@@ 3.1 Original 3DGS Representation
To make optimization stable, i.e. guarantee that $\bm{\Sigma}$ is positive semi-definite, the covariance matrix is further decomposed into rotation $R$ and scaling $S$:
@@
Color is given as spherical harmonics (SH) of degree $k$ and is thus view-dependent.
The Gaussian parameters are optimized via a differentiable rasterizer, that projects the current 3D Gaussians to the image space and compares against the ground truth images. This rasterizer relies on an efficient algorithm for sorting the Gaussians and tiling the image space, which leads to very fast training and rendering. Alongside the optimizations, 3DGS employs an adaptive mechanism for pruning and densification, to make sure that the set of gaussians represents the space effectively. For more details, we refer the reader to the 3DGS paper [20].